A (B+25.0) g mass is hung on a spring. As a result the spring stretches (8.50+A) cm. If the object is then pulled an additional 3.00 cm downward and released, what is the period of the resulting oscillation? Give your answer in seconds with 3 significant figures.

A= 10

B=910
Answer #1


given
m = 910 + 25

= 935 g

= 0.935 kg

x = 8.5 + 10

= 18.5 cm

= 0.185 m

let k is the spring constant.

in the equilibrium, F_spring = Fg

k*x = m*g

k = m*g/x

= 0.935*9.8/0.185

= 49.5 N/m

time period of oscillations,

T = 2*pi*sqrt(m/k)

= 2*pi*sqrt(0.935/49.5)

= 0.863 s <<<<<<<<<-------------Answer